two fixed charges 1.07 micro coulumb and -3.28 micro coulomb are 61.8 cm apart.where may a third charge be located so that no net force acts on it?

Grade:12th pass

1 Answers

Vikas TU
5 years ago
let q be the charge and is placed at x distance from left of the 1.07 uC charge.
Net force acting on it would be=>
K(1.07)q/x^2 – K(3.28)q/(x + 0.618)^2 = 0
1.07/x^2 = 3.28/(x + 0.618)^2
(x + 0.618)^2/x^2 = 3.28/1.07 => 3.065
(x + 0.618)/x = 1.75
(x + 0.618) = 1.75x
0.75x = 0.618
x = 0.618/0.75 = 82.3 cm.